# The Dynamical History of the Kepler-221 Planet System

### v1

@@
Kepler-221 is a G-type star hosting four planets. In this system, planets b, c, and e are in (or near) a 6:3:1 three-body resonance even though the planets’ period ratios show significant departures from exact two-body commensurability. Importantly, the intermediate planet d is not part of the resonance chain. To reach this resonance configuration, we propose a scenario in which there were originally five planets in the system in a chain of first-order resonances. After disk dispersal, the resonance chain became unstable and two planets quickly merged to become the current planet d. In addition, the b/c/e three-body resonance was re-established. We run N-body simulations using REBOUND to investigate the parameter space under which this scenario can operate. We find that our envisioned scenario is possible when certain conditions are met. First, the reformation of the three-body resonance after planet merging requires convergent migration between planets b and c. Second, as has previously pointed out, an efficient damping mechanism must operate to power the expansion of the b/c/e system. We find that planet d plays a crucial role during the orbital expansion phase due to destabilizing encounters of a three-body resonance between c, d, and e. A successful orbital expansion phase puts constraints on the planet properties in the Kepler-221 system including the planet mass ratios and the tidal quality factors for the planets. Our model can also be applied to other planet systems in resonance, such as Kepler-402 and K2-138.
@@ 1 Introduction
In the past two decades, with various transit and radial-velocity surveys, the number of multi-exoplanetary systems detected has exceeded 900111https://exoplanetarchive.ipac.caltech.edu. Among these, there is an excess of systems with adjacent planets in period ratios close to integer (Fabrycky et al., 2014; Steffen & Hwang, 2015; Huang & Ormel, 2023; Hamer & Schlaufman, 2024; Dai et al., 2024), which are often identified with mean-motion resonances (MMRs). MMRs in these multi-planetary systems are believed to form early, in the gas-rich disk phase, through Type-I convergent migration (Goldreich & Tremaine, 1979; Lin & Papaloizou, 1979), with the exact resonant state related to the disk and planet properties (Ogihara & Kobayashi, 2013; Kajtazi et al., 2023; Huang & Ormel, 2023; Batygin & Petit, 2023). Therefore, resonant systems preserve relics of formation processes and unveil the footprint of the dynamical history of multi-planetary systems (Snellgrove et al., 2001; Papaloizou & Szuszkiewicz, 2005; Teyssandier & Libert, 2020; Huang & Ormel, 2022; Pichierri et al., 2024).
The formal dynamical identification of a two-body resonance (2BR) is in terms of an angle (Murray & Dermott, 1999), which relies on a precise knowledge of the argument of pericenter and is observationally hard to assess. However, it is much easier to recognize three planets in three-body resonance (3BR) because the corresponding resonance angle only depends on the angular positions of the planets (their mean longitudes). Multiple exoplanetary systems have been confirmed to host planets in 3BRs, including GJ 876 (Rivera et al., 2010), Kepler-80 (Xie, 2013; MacDonald et al., 2016, 2021), Kepler-221 (Rowe et al., 2014; Goldberg & Batygin, 2021), Kepler-60 (Goździewski et al., 2016), Kepler-223 (Mills et al., 2016), TRAPPIST-1 (Gillon et al., 2017; Luger et al., 2017; Agol et al., 2021; Huang & Ormel, 2022), K2-138 (Christiansen et al., 2018; Leleu et al., 2019, 2021a; Cerioni & Beaugé, 2023), HD-158259 (Hara et al., 2020), TOI-178 (Leleu et al., 2021b), TOI-1136 (Dai et al., 2023), HD-110067 (Luque et al., 2023; Lammers & Winn, 2024) and Kepler-402 (see Section 6.4).
@@
Among the planetary systems with 3BR, Kepler-221 is unique because it hosts a non-adjacent 3BR containing a second-order 2BR (Goldberg & Batygin, 2021). Kepler-221 is a G5-type star ($R_{s}=0.82R_{\odot}$) hosting four planets (b, c, d, and e) at a distance of 385 $\mathrm{pc}$, with the system information listed in Table 1 (Rowe et al., 2014; Berger et al., 2018). Kepler-221 system is likely to be relatively young, with its large lithium abundance suggesting that the system is younger than the Hyades with an age estimated around 650 Myr (Berger et al., 2018; Goldberg & Batygin, 2021). In the Kepler-221 system, the radii of the planets are well constrained but the masses of the planets are unknown due to weak TTVs (Berger et al., 2018; Goldberg & Batygin, 2021). In the system, planets b, c, and e are in a 6:3:1 3BR with the period ratios relatively far away from the integer period ratio (the period ratio between planets b and c is 2.035 and the period ratio between c and e is 3.228). However, planet d is not in any resonance with other planets.
@@
If the resonance chain in the Kepler-221 planetary system formed due to convergent migration during the disk phase, it is generally expected that the planets will be locked into a 1st-order resonance chain (Lee & Peale, 2002). However, except for planets b and c, this is not what is observed in the Kepler-221 system. Two observations stand out. First, planets c and e are close to (or in) 3:1 resonance, which is weaker than a 1st-order resonance. Second, planet d is clearly not in resonance with other planets (not even close to any). Both these two features cannot be explained by a simple disk migration model.
@@
To explain the unique dynamical configuration of the Kepler-221 system, we propose a new dynamical model in which there were originally five planets in a 1st-order resonance chain, as would be expected from migration theory. The resonance breaks in the post-disk phase due to dynamical instability from infalling materials from a potential debris disk (Izidoro et al., 2017, 2021; Liu et al., 2022; Nagpal et al., 2024) or planetesimal scattering (Raymond et al., 2022; Ghosh & Chatterjee, 2023; Griveaud et al., 2024; Wu et al., 2024). Therefore, the system became unstable and two planets merged into the current planet d while the three-body resonance between the other three planets (b, c, and e) reformed. Over long timescales, tidal dissipation fueled the expansion of these planets to their current period ratios. To investigate this scenario quantitatively, a modular approach is adopted, where we investigate the viability of each of these steps separately. The advantage of this approach is that the resulting conclusions are self-contained. In particular, when planet b/c/e 3BR expands toward the observed period ratios, we find that planet d puts additional constraints on the planet’s tidal parameters and masses, which are independent of the detailed formation scenario.
@@
This paper is structured as follows. In Section 2, we present the dynamical configuration of the Kepler-221 system and the new four-phased dynamical model. In Section 3, we present a consistent simulation throughout all phases which successfully reproduced the observation configuration. Then we use N-body simulations to verify our model and conduct parameter studies following a chronological order. In Section 4, the evolution of the system up to the point of three-body reformation is discussed. Section 5 investigates the orbital expansion phase, which puts constraints on the mass ratio of planets in resonance. According to the simulation results, we assess the five-planet formation model in Section 6. Finally, we summarize our results and present them in Section 7.
@@ 2.1 Resonance in the Kepler-221 Planetary System
$\phi_{1,2,X}=(j+o)l_{1}-jl_{2}-o\varpi_{X}$ (1)
where $l_{1}$ and $l_{2}$ stand for the mean longitude for the inner and outer planet, $\varpi_{X}$ stands for either the longitude of periapsis for planet 1 ($X=1$) or planet 2 ($X=2$), which, respectively, correspond to the inner and outer resonance angle. In Equation (1) $j$ and $o$ are integers called the resonance number and resonance order, respectively. For two-body resonances, 1st-order resonances are more stable ($o=1$) than higher-order resonances ($o>1$) (Murray & Dermott, 1999). For two planets in resonance, either one or both inner and outer resonance angles should librate about a fixed value. For planets not in resonance, the resonance angle instead circulates from 0 to $2\pi$.
@@
$\phi_{3BR}=p\lambda_{1}-r\lambda_{2}+q\lambda_{3}$ (2)
where $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are the mean longitude for the three planets. The quantity $(p+q-r)$ is called the 3BR order. Planets are locked in 3BR if such an angle librates around a fixed value. 0th-order 3BR and 1st-order 3BR are generally stable and hard to break (Petit, 2021). For the Kepler-221 system, the period ratio between planets b and c is 2.035 and the period ratio between c and e is 3.228, which are close to 2:1 and 3:1, respectively. For simplicity, a librating $\phi_{3BR}=p\lambda_{1}-r\lambda_{2}+q\lambda_{3}$ will be referred to using the $(p,q,r)$ 3BR notation in the following. Therefore, planets b, c, and e might be in the $(2,3,5)$ 0th-order 3BR. The B-value of 3BR is defined by taking the time derivative of Equation (2):
$B=\dot{\phi}_{3BR}=pn_{1}-rn_{2}+qn_{3}$ (3)
where $n_{1}$, $n_{2}$ and $n_{3}$ are the mean motion for the three planets and $B$ is the B-value. For planets in 3BR, because their 3BR angle liberates around a certain value, the B-value averaged over time should be close to 0.
@@
$B_{\mathrm{norm}}=|B|/\langle n\rangle.$ (4)
where $\langle n\rangle$ is the average of the mean motion of the planets. For planet b, c, and e of Kepler-221 $B_{\mathrm{norm}}=1.03\times 10^{-5}$ when $(p,q,r)=(2,3,5)$. Compared with other confirmed 3BR pairs such as K2-138 (Christiansen et al., 2018), Kepler-80 (MacDonald et al., 2016), and Kepler-60 (Goździewski et al., 2016), the normalized B-value for Kepler-221 is statistically closer to 0, which implies they are in $(2,3,5)$ 0th-order 3BR (Goldberg & Batygin, 2021).
@@ 2.2 The Formation of the Kepler-221 Planetary System
Figure 1: Schematic of the formation model for the Kepler-221 planet system investigated in this study. Sequential migration traps five planets in a chain of 1st-order resonances (panels a+b). After disk dispersal, the resonance chain breaks (panel c) triggering a dynamical instability that results in the merger of planets $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ (panel d). Convergent migration between planets b and c re-establishes the b, c, and e resonance chain (panel e). On evolutionary timescales (${\sim}1\,\mathrm{Gyr}$) the b, c, and e resonance chain expands to the present-day period ratios due to tidal dissipation (panel f).
To solve the puzzle of the formation of the b/c/e three-body resonance chain, we hypothesize that Kepler-221 originally harbored five planets, but two of them ($\mathrm{d}_{1}$ and $\mathrm{d}_{2}$) merged to form planet d. The model that we envision consists of four phases, as illustrated in Figure 1. In chronological order:
@@
    In the disk phase, five planets formed and then migrated inward due to Type-I migration. All planets are assumed to stay beyond the inner edge (Liu et al., 2017). This leads to convergent migration and the adjacent planets are locked into a stable 1st-order two-body resonance chain: 2:1, 3:2, 4:3, and 3:2. The disk then disperses and the stable resonance chain remains, as shown in panel b in Figure 1.
@@
    In the third phase – the post-collision phase – the b/c/e 3BR reforms (panels d and e of Figure 1). Planets b, c, and e are initially not in resonance after the collision, but they stay very close to the 0th-order 6:3:1 resonance position. Tidal damping in the system will damp the planets’ eccentricities. Convergent migration due to tidal damping results in the reformation of the b, c, and e 3BR.
@@
    After reforming the 0th-order 3BR, the system expands to the current period ratio – the orbital expansion phase (see panel f of Figure 1). In this last phase, tidal damping will continue to expand the system while preserving the 3-body resonance (Goldberg & Batygin, 2021). This is because tidal damping preserves the angular momentum but dissipates the energy. As a result, the period ratio between planets in resonance in the system expands and the system reaches the observed position at present.
In this scenario, the first three phases are essential to explain the presence of the 3-body b/c/e resonance chain in Kepler-221. The last phase explains the departure of period ratios from exact integer ratios through b/c/e 3BR expansion. This phase is independent of the details of the previous three phases; it only requires that planets b, c, and e are in resonance.
@@ 2.3 Numerical Implementation
$\bm{F}_{\mathrm{disk,i}}=-\frac{\mathbf{v}_{\mathrm{i}}}{2t_{\mathrm{a,i}}}-\frac{2\left(\mathbf{v}_{\mathrm{i}}\cdot\mathbf{r}_{\mathrm{i}}\right)\mathbf{v}_{\mathrm{i}}}{\left|r_{\mathrm{i}}\right|^{2}t_{\mathrm{e,i}}}$ (5)
@@ 3 Charting the Course: A Successful Simulation
Table 3: Parameter setups in different simulation stage corresponding to figures from Figure 2 to Figure 6.
@@
444aHere the parameters denote – $m_{\mathrm{d2}}/m_{\mathrm{d1}}$: the mass ratio between planet d2 and d1 before merging, $t_{\mathrm{a,I}}$: semi-major axis damping timescale in the disk phase, $t_{\mathrm{a,I}}/t_{\mathrm{e,I}}$: the ratio between semi-major axis damping and eccentricity damping timescale in disk phase, $t_{\mathrm{d}}$: the disk dispersal timescale, $\Delta a/a$: the instantaneous kick applied to the semi-major axis of planet d2 to break the resonances, $Q_{\mathrm{phy,x}}$: the effective tidal damping parameter for planet $x$ (see Section 2.3), $t^{b}_{\mathrm{a0,III}},t^{b}_{\mathrm{a0,IV}}$: the equivalent semi-major axis migration timescale due to torque on planet b in Phase III and Phase IV, $t^{b}_{\mathrm{\Gamma,III}},t^{b}_{\mathrm{\Gamma,IV}}$: the exponential decay timescale of the torque on planet b in Phase III and Phase IV, $P_{\mathrm{d}}/P_{\mathrm{c}}$: the period ratio of planet c and d after planet merging. Let’s use $t_{a0}$ for the initial damping timescale and $t_{\Gamma}$ for the timescale on which the torque decays, instead.
@@
After disk dispersal (end of Phase I), the planets in the simulation stay close to the resonance position (integer period ratio), which is typical for young resonant planetary systems (Hamer & Schlaufman, 2024; Dai et al., 2024). Due to the adopted damping parameters the the planets end up with a relatively high eccentricity of around 0.01-0.1 (Teyssandier & Terquem, 2014; Papaloizou et al., 2018; Yang & Li, 2024).
@@
Panels c and g represent the post-collision phase in Figure 2. After the merging of $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$, the eccentricity of planet d decreases because planets $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ merged at their respective perihelion and aphelion distances, leading to a more circular orbit for planet d after merging (Kokubo & Ida, 1995). Specifically, the eccentricity of planet $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ is around 0.1 before the merger while the eccentricity of planet d is around 0.05 after merging. Tidal damping also continuously decreases the eccentricity. In the post-collision phase, planets b and c undergo convergent migration due to the outward torque that is applied to planet b (see Section 5.2.1). Such an outward torque on planet b is essential to reform the resonance chain (see Section 4.3). The 2:1 two-body resonance between planet b and c reforms first (after $1.7\,\mathrm{Myr}$; the two-body angle is not shown in Figure 2). Thereafter, the c and e two-body 3:1 resonance and the b, c, and e 3BR reform simultaneously at around $5.7\,\mathrm{Myr}$. The conditions are now in place for the planet period ratios $P_{\mathrm{c}}/P_{\mathrm{b}}$ and $P_{\mathrm{e}}/P_{\mathrm{c}}$ to expand along the 0th-order 3BR line (blue-dashed line in Figure 2.
@@
Panel d and h in Figure 2 show the evolution of the b/c/e and c/d/e period ratio during the orbital expansion phase (Phase IV). After the b/c/e 3BR reforms, tidal damping operates to expand the period ratio of the planets in resonance, migrating planets b and c inward and planet e outward. Initially, the planets are close to their corresponding 2:1 and 3:1 period ratios. The dashed lines in Figure 2 indicate different 3BR characterized by $p$, $q$ and $r$ in Equation (3):
$\frac{P_{\mathrm{3}}}{P_{\mathrm{2}}}=\left(\frac{r}{q}-\frac{pP_{\mathrm{2}}}{qP_{\mathrm{1}}}\right)^{-1}$ (7)
@@
where $P_{\mathrm{1}}$, $P_{\mathrm{2}}$, $P_{\mathrm{3}}$ correspond to the inner, middle, and outer planets (i.e., b/c/e in Figure 2d and c/d/e in Figure 2h). In panel d of Figure 2, tidal dissipation moves the planets along the blue dashed line corresponding to the $(2,3,5)$ 0th-order 3BR between b, c, and e (Charalambous et al., 2018; Papaloizou et al., 2018). At the same time, the period ratio of planet c/d/e also evolves along a line that avoids interaction with certain c/d/e 3BRs as shown in Figure 2h. The slope of the expansion seen in Figure 2h is related to the masses of the planets, as will be discussed in Section 5.1. After about $2\,\mathrm{Gyr}$ of expansion, the period ratios of the planets are consistent with their observed values, implying a successful reproduction of the dynamical configuration of the Kepler-221 system.

### v2

@@
Kepler-221 is a G-type star hosting four planets. In this system, planets b, c, and e are in (or near) a 6:3:1 three-body resonance even though the planets’ period ratios show significant departures from exact two-body commensurability. Importantly, the intermediate planet d is not part of the resonance chain. To reach this resonance configuration, we propose a scenario in which there were originally five planets in the system in a chain of first-order resonances. After disk dispersal, the resonance chain became unstable, and two planets quickly merged to become the current planet d. In addition, the (b, c, e) three-body resonance was re-established. We ran N-body simulations using REBOUND to investigate the parameter space under which this scenario can operate. We find that our envisioned scenario is possible when certain conditions are met. First, the reformation of the three-body resonance after planet merging requires convergent migration between planets b and c. Second, as has been previously pointed out, an efficient damping mechanism must operate to power the expansion of the (b, c, e) system. We find that planet d plays a crucial role during the orbital expansion phase due to destabilizing encounters of a three-body resonance between c, d, and e. A successful orbital expansion phase puts constraints on the planet properties in the Kepler-221 system including the planet mass ratios and the tidal quality factors for the planets. Our model can also be applied to other planet systems in resonance, such as Kepler-402 and K2-138.
@@ 1 Introduction
In the past two decades, with various transit and radial-velocity surveys, the number of multi-exoplanetary systems detected has exceeded 900.111https://exoplanetarchive.ipac.caltech.edu Among these, there is an excess of systems with adjacent planets in period ratios close to integer (Fabrycky et al., 2014; Steffen & Hwang, 2015; Huang & Ormel, 2023; Hamer & Schlaufman, 2024; Dai et al., 2024), which are often identified with mean-motion resonances (MMRs). The MMRs in these multi-planetary systems are believed to form early, in the gas-rich disk phase, through Type-I convergent migration (Goldreich & Tremaine, 1979; Lin & Papaloizou, 1979), with the exact resonant state related to the disk and planet properties (Ogihara & Kobayashi, 2013; Kajtazi et al., 2023; Huang & Ormel, 2023; Batygin & Petit, 2023). Therefore, resonant systems preserve relics of formation processes and can unveil the footprint of the dynamical history of multi-planetary systems (Snellgrove et al., 2001; Papaloizou & Szuszkiewicz, 2005; Teyssandier & Libert, 2020; Huang & Ormel, 2022; Pichierri et al., 2024).
The formal dynamical identification of a two-body resonance (2BR) is in terms of an angle (Murray & Dermott, 1999), which relies on a precise knowledge of the argument of pericenter and is observationally hard to assess. However, it is much easier to recognize three planets in a three-body resonance (3BR) because the corresponding resonance angle only depends on the angular positions of the planets (their mean longitudes). Multiple exoplanetary systems have been confirmed to host planets in 3BRs, including GJ 876 (Rivera et al., 2010), Kepler-80 (Xie, 2013; MacDonald et al., 2016, 2021), Kepler-221 (Rowe et al., 2014; Goldberg & Batygin, 2021), Kepler-60 (Goździewski et al., 2016), Kepler-223 (Mills et al., 2016), TRAPPIST-1 (Gillon et al., 2017; Luger et al., 2017; Agol et al., 2021; Huang & Ormel, 2022), K2-138 (Christiansen et al., 2018; Leleu et al., 2019, 2021a; Cerioni & Beaugé, 2023), HD-158259 (Hara et al., 2020), TOI-178 (Leleu et al., 2021b), TOI-1136 (Dai et al., 2023), HD-110067 (Luque et al., 2023; Lammers & Winn, 2024), and Kepler-402 (see Section 6.4).
@@
Among the planetary systems with a 3BR, Kepler-221 is unique because it hosts a non-adjacent 3BR containing a second-order 2BR (Goldberg & Batygin, 2021). Kepler-221 is a G5-type star ($R_{s}=0.82R_{\odot}$) hosting four planets (b, c, d, and e) at a distance of 385 $\mathrm{pc}$, with the system information listed in Table 1 (Rowe et al., 2014; Berger et al., 2018). Kepler-221 system is likely to be relatively young, with its large lithium abundance suggesting that the system is younger than the Hyades with an age estimated around 650 Myr (Berger et al., 2018; Goldberg & Batygin, 2021). In the Kepler-221 system, the radii of the planets are well constrained, but the masses of the planets are unknown due to weak TTVs (Berger et al., 2018; Goldberg & Batygin, 2021). In the system, planets b, c, and e are in a 6:3:1 3BR with the period ratios relatively far away from the integer period ratio (the period ratio between planets b and c is 2.035, and the period ratio between c and e is 3.228). However, planet d is not in any resonance with other planets.
@@
If the resonance chain in the Kepler-221 planetary system formed due to convergent migration during the disk phase, it is generally expected that the planets will be locked into a first-order resonance chain (Lee & Peale, 2002). However, except for planets b and c, this is not what is observed in the Kepler-221 system. Two observations stand out. First, planets c and e are close to (or in) 3:1 resonance, which is weaker than a first-order resonance. Second, planet d is clearly not in resonance with other planets (not even close to any). These two features cannot be explained by a simple disk migration model.
@@
To explain the unique dynamical configuration of the Kepler-221 system, we propose a new dynamical model in which there were originally five planets in a first-order resonance chain, as would be expected from migration theory. The resonance breaks in the post-disk phase due to dynamical instability from infalling materials from a potential debris disk (Izidoro et al., 2017, 2021; Liu et al., 2022; Nagpal et al., 2024) or planetesimal scattering (Raymond et al., 2022; Ghosh & Chatterjee, 2023; Griveaud et al., 2024; Wu et al., 2024). Therefore, the system became unstable, and two planets merged into the current planet d, while the three-body resonance between the other three planets (b, c, and e) reformed. Over long timescales, tidal dissipation fueled the expansion of these planets to their current period ratios. To investigate this scenario quantitatively, a modular approach was adopted, where we investigated the viability of each of these steps separately. The advantage of this approach is that the resulting conclusions are self-contained. In particular, when planet (b, c, e) 3BR expands toward the observed period ratios, we find that planet d puts additional constraints on the planet’s tidal parameters and masses, which are independent of the detailed formation scenario.
@@
This paper is structured as follows. In Section 2, we present the dynamical configuration of the Kepler-221 system and the new four-phased dynamical model. In Section 3, we present a consistent simulation throughout all phases that successfully reproduced the observation configuration. Then we use N-body simulations to verify our model and conduct parameter studies following a chronological order. In Section 4, the evolution of the system up to the point of three-body reformation is discussed. Section 5 investigates the orbital expansion phase, which puts constraints on the mass ratio of planets in resonance. According to the simulation results, we assess the five-planet formation model in Section 6. Finally, we summarize our results and present them in Section 7.
@@ 2.1 Resonance in the Kepler-221 planetary system
$\phi_{1,2,X}=(j+o)l_{1}-jl_{2}-o\varpi_{X},$ (1)
where $l_{1}$ and $l_{2}$ stand for the mean longitude for the inner and outer planet, $\varpi_{X}$ stands for either the longitude of periapsis for planet 1 ($X=1$) or planet 2 ($X=2$), which, respectively, correspond to the inner and outer resonance angle. In Equation (1), $j$ and $o$ are integers called the resonance number and resonance order, respectively. For two-body resonances, first-order resonances are more stable ($o=1$) than higher-order resonances ($o>1$) (Murray & Dermott, 1999). For two planets in resonance, either one or both inner and outer resonance angles should librate about a fixed value. For planets not in resonance, the resonance angle instead circulates from 0 to $2\pi$.
@@
$\phi_{3BR}=p\lambda_{1}-r\lambda_{2}+q\lambda_{3},$ (2)
where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are the mean longitude for the three planets. The quantity $(p+q-r)$ is called the 3BR order. Planets are locked in 3BR if such an angle librates around a fixed value. The zeroth-order 3BR and first-order 3BR are generally stable and hard to break (Petit, 2021). For the Kepler-221 system, the period ratio between planets b and c is 2.035, and the period ratio between c and e is 3.228, which is close to 2:1 and 3:1, respectively. For simplicity, a librating $\phi_{3BR}=p\lambda_{1}-r\lambda_{2}+q\lambda_{3}$ is referred to by using the $(p,q,r)$ 3BR notation in the following. Therefore, planets b, c, and e might be in the $(2,3,5)$ zeroth-order 3BR. The B-value of a 3BR is defined by taking the time derivative of Equation (2):
$B=\dot{\phi}_{3BR}=pn_{1}-rn_{2}+qn_{3},$ (3)
where $n_{1}$, $n_{2}$, and $n_{3}$ are the mean motion for the three planets and $B$ is the B-value. For planets in 3BR, because their 3BR angle librates around a certain value, the B-value averaged over time should be close to zero.
@@
$B_{\mathrm{norm}}=|B|/\langle n\rangle,$ (4)
where $\langle n\rangle$ is the average of the mean motion of the planets. For planets b, c, and e of Kepler-221, $B_{\mathrm{norm}}=1.03\times 10^{-5}$ when $(p,q,r)=(2,3,5)$. Compared with other confirmed 3BR pairs such as K2-138 (Christiansen et al., 2018), Kepler-80 (MacDonald et al., 2016), and Kepler-60 (Goździewski et al., 2016), the normalized B-value for Kepler-221 is statistically closer to zero, which implies that planets b, c, and e are in $(2,3,5)$ zeroth-order 3BR (Goldberg & Batygin, 2021).
@@ 2.2 The Formation of the Kepler-221 planetary system
Figure 1: Schematic of the formation model for the Kepler-221 planet system investigated in this study. Sequential migration traps five planets in a chain of first-order resonances (panels a+b). After disk dispersal, the resonance chain breaks (panel c) triggering a dynamical instability that results in the merger of planets $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ (panel d). Convergent migration between planets b and c re-establishes the b, c, and e resonance chain (panel e). On evolutionary timescales (${\sim}1\,\mathrm{Gyr}$) the b, c, and e resonance chain expands to the present-day period ratios due to tidal dissipation (panel f).
To solve the puzzle of the formation of the (b, c, e) three-body resonance chain, we hypothesize that Kepler-221 originally harbored five planets, but two of them ($\mathrm{d}_{1}$ and $\mathrm{d}_{2}$) merged to form planet d. The model that we envision consists of four phases, as illustrated in Figure 1. In chronological order:
@@
    In the disk phase, five planets formed and then migrated inward due to Type-I migration. All planets are assumed to stay beyond the inner edge (Liu et al., 2017). This leads to convergent migration and the adjacent planets are locked into a stable first-order two-body resonance chain: 2:1, 3:2, 4:3, and 3:2. The disk then disperses and the stable resonance chain remains, as shown in panel b in Figure 1.
@@
    In the third phase – the post-collision phase – the (b, c, e) 3BR reforms (panels d and e of Figure 1). Planets b, c, and e are initially not in resonance after the collision, but they stay very close to the zeroth-order 6:3:1 resonance position. Tidal damping in the system will damp the planets’ eccentricities. Convergent migration due to tidal damping results in the reformation of the b, c, and e 3BR.
@@
    After reforming the zeroth-order 3BR, the system expands to the current period ratio – the orbital expansion phase (see panel f of Figure 1). In this last phase, tidal damping will continue to expand the system while preserving the 3-body resonance (Goldberg & Batygin, 2021). This is because tidal damping preserves the angular momentum but dissipates the energy. As a result, the period ratio between planets in resonance in the system expands and the system reaches the observed position at present.
In this scenario, the first three phases are essential to explain the presence of the 3-body (b, c, e) resonance chain in Kepler-221. The last phase explains the departure of period ratios from exact integer ratios through (b, c, e) 3BR expansion. This phase is independent of the details of the previous three phases; it only requires that planets b, c, and e are in resonance.
@@ 2.3 Numerical implementation
$\bm{F}_{\mathrm{disk,i}}=-\frac{\mathbf{v}_{\mathrm{i}}}{2t_{\mathrm{a,i}}}-\frac{2\left(\mathbf{v}_{\mathrm{i}}\cdot\mathbf{r}_{\mathrm{i}}\right)\mathbf{v}_{\mathrm{i}}}{\left|r_{\mathrm{i}}\right|^{2}t_{\mathrm{e,i}}},$ (5)
@@ 3 Charting the course: a successful simulation
Table 3: Parameter setups in different simulation stages corresponding to figures from Figure 2 to Figure 6.
@@
444aHere the parameters denote – $m_{\mathrm{d2}}/m_{\mathrm{d1}}$: the mass ratio between planet d2 and d1 before merging, $t_{\mathrm{a,I}}$: semi-major axis damping timescale in the disk phase, $t_{\mathrm{a,I}}/t_{\mathrm{e,I}}$: the ratio between semi-major axis damping and eccentricity damping timescale in disk phase, $t_{\mathrm{d}}$: the disk dispersal timescale, $\Delta a/a$: the instantaneous kick applied to the semi-major axis of planet d2 to break the resonances, $Q_{\mathrm{phy,x}}$: the effective tidal damping parameter for planet $x$ (see Section 2.3), $t^{b}_{\mathrm{a0,III}},t^{b}_{\mathrm{a0,IV}}$: the equivalent semi-major axis migration timescale due to torque on planet b in Phase III and Phase IV, $t^{b}_{\mathrm{\Gamma,III}},t^{b}_{\mathrm{\Gamma,IV}}$: the exponential decay timescale of the torque on planet b in Phase III and Phase IV, $P_{\mathrm{d}}/P_{\mathrm{c}}$: the period ratio of planet c and d after planet merging.
@@
After disk dispersal (end of Phase I), the planets in the simulation stay close to the resonance position (integer period ratio), which is typical for young resonant planetary systems (Hamer & Schlaufman, 2024; Dai et al., 2024). Due to the adopted damping parameters the planets end up with a relatively high eccentricity of around 0.01-0.1 (Teyssandier & Terquem, 2014; Papaloizou et al., 2018; Yang & Li, 2024).
@@
Panels c and g represent the post-collision phase in Figure 2. After the merging of $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$, the eccentricity of planet d decreases because planets $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ merged at their respective perihelion and aphelion distances, leading to a more circular orbit for planet d after merging (Kokubo & Ida, 1995). Specifically, the eccentricity of planet $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ is around 0.1 before the merger while the eccentricity of planet d is around 0.05 after merging. Tidal damping also continuously decreases the eccentricity. In the post-collision phase, planets b and c undergo convergent migration due to the outward torque that is applied to planet b (see Section 5.2.1). Such an outward torque on planet b is essential to reform the resonance chain (see Section 4.3). The 2:1 two-body resonance between planet b and c reforms first (after $1.7\,\mathrm{Myr}$; the two-body angle is not shown in Figure 2). Thereafter, the c and e two-body 3:1 resonance and the b, c, and e 3BR reform simultaneously at around $5.7\,\mathrm{Myr}$. The conditions are now in place for the planet period ratios $P_{\mathrm{c}}/P_{\mathrm{b}}$ and $P_{\mathrm{e}}/P_{\mathrm{c}}$ to expand along the zeroth-order 3BR line (blue-dashed line in Figure 2.
@@
Panel d and h in Figure 2 show the evolution of the (b, c, e) and (c, d, e) period ratios during the orbital expansion phase (Phase IV). After the (b, c, e) 3BR reforms, tidal damping operates to expand the period ratio of the planets in resonance, migrating planets b and c inward and planet e outward. Initially, the planets are close to their corresponding 2:1 and 3:1 period ratios. The dashed lines in Figure 2 indicate different 3BRs characterized by $p$, $q$, and $r$ in Equation (3):
$\frac{P_{\mathrm{3}}}{P_{\mathrm{2}}}=\left(\frac{r}{q}-\frac{pP_{\mathrm{2}}}{qP_{\mathrm{1}}}\right)^{-1},$ (7)
@@
where $P_{\mathrm{1}}$, $P_{\mathrm{2}}$, $P_{\mathrm{3}}$ correspond to the inner, middle, and outer planets (i.e., (b, c, e) in Figure 2d and (c, d, e) in Figure 2h). In panel d of Figure 2, tidal dissipation moves the planets along the blue dashed line corresponding to the $(2,3,5)$ zeroth-order 3BR between b, c, and e (Charalambous et al., 2018; Papaloizou et al., 2018). At the same time, the period ratio of planets c, d, and e also evolves along a line that avoids interaction with certain (c, d, e) 3BRs as shown in Figure 2h. The slope of the expansion seen in Figure 2h is related to the masses of the planets, as is discussed in Section 5.1. After about $2\,\mathrm{Gyr}$ of expansion, the period ratios of the planets are consistent with their observed values, implying a successful reproduction of the dynamical configuration of the Kepler-221 system.